# Language-Driven Closed-Loop Grasping with Model-Predictive Trajectory Replanning

### v1

@@ IV-A Statistical results of vision modules
Considering the same test setup depicted in Figure 3 and all ten objects shown in Figure 4, we conducted statistical analysis on the inference speeds and success rates of using another state-of-the-art language-driven vision module, namely CNOS [26] with FastSAM [25], in conjunction with the presented object localization module. For all test cases, the prompt ”Please grasp the [object]” was used, with the placeholder [object] being replaced by terms such as ”orange drill,” ”blue block,” ”orange block,” ”green block,” etc. Each object underwent five trials. Table II presents the computation times of the inference processes of the vision modules using different GPUs. The term ”oFP & LKF” refers to the online version of FoundationPose with a linear Kalman filter, which is utilized in our object pose localization module. It is noteworthy that the CNOS & FastSAM models are pretty slow, requiring several seconds to segment objects through language commands. Conversely, the lightweight OWLv2 requires approximately to compute a binary mask of the segmented object via a prompt. It is not surprising that with a Nvidia RTX 4090, an update rate of approximately can be achieved using the language-driven module employing OWLv2 [6], and approximately with the proposed object localization module. Under identical lighting conditions and prompts across 50 trials, OWLv2 detected objects in 46 instances, whereas the other methods succeeded in only 40 cases. It is important to note that once an object is detected and the binary mask is provided, the object localization model can reliably track the object.
@@ IV-C Robotic experiments
AnyGrasp [35], which does not utilize natural language prompts, achieved 37 successful trials out of 50 without prompts. The LGD [36] approach encountered 27 failures due to incorrect object detections in cluttered scenarios. Unlike LGD [36] and AnyGrasp [35], both VoxPoser [24] and our modular framework consider the pre-grasp pose for all cases, as shown in Figure 3.
@@ V Conclusions
We have introduced a novel zero-shot modular framework designed to address the challenges of language-driven manipulation of dynamic objects within a closed-loop control system. Our framework leverages a vision-language model for natural language processing and real-time object segmentation, allowing for the seamless integration of 6D object pose localization and receding-horizon trajectory optimization to achieve precise and smooth robotic movements. The intensive experimental results highlight the robustness and efficiency of our framework, achieving a 92% success rate and superior inference speeds compared to state-of-the-art methods. Furthermore, the object localization module demonstrated high reliability with minimal positional and orientation errors, particularly in dynamic environments requiring real-time replanning. Future work will focus on further improving 6D pose estimation by continuously localizing object segmentation to prevent tracking loss and detecting affordance grasping objects through a language-driven approach.

### v2

@@ IV-A Statistical results of vision modules
Considering the same test setup depicted in Figure 3 and all ten objects shown in Figure 4, we conducted statistical analysis on the inference speeds and success rates of using another state-of-the-art language-driven vision module, namely CNOS [25] with FastSAM [24], in conjunction with the presented object localization module. For all test cases, the prompt ”Please grasp the [object]” was used, with the placeholder [object] being replaced by terms such as ”orange drill,” ”blue block,” ”orange block,” ”green block,” etc. Each object underwent five trials. Table II presents the computation times of the inference processes of the vision modules using different GPUs. The term ”oFP & KF” refers to the online version of FoundationPose with a Kalman filter, which is utilized in our object pose localization module. It is noteworthy that the CNOS & FastSAM models are pretty slow, requiring several seconds to segment objects through language commands. Conversely, the lightweight OWLv2 requires approximately to compute a binary mask of the segmented object via a prompt. It is not surprising that with a Nvidia RTX 4090, an update rate of approximately can be achieved using the language-driven module employing OWLv2 [5], and approximately with the proposed object localization module. Under identical lighting conditions and prompts across 50 trials, OWLv2 detected objects in 46 instances, whereas the other methods succeeded in only 40 cases. It is important to note that once an object is detected and the binary mask is provided, the object localization model can reliably track the object.
@@ IV-C Robotic experiments
AnyGrasp [33], which does not utilize natural language prompts, achieved 37 successful trials out of 50 without prompts. The LGD [34] approach encountered 27 failures due to incorrect object detections in cluttered scenarios. Unlike LGD [34] and AnyGrasp [33], both VoxPoser [23] and our modular framework consider the pre-grasp pose for all cases, as shown in Figure 3. Note that, once the grasping pose is determined using AnyGrasp and LGD, the robot utilizes a standard inverse kinematic algorithm [29] alongside a joint space controller to guide its movements toward the object.
@@ V Conclusions
We have introduced a novel zero-shot modular framework designed to address the challenges of language-driven manipulation of dynamic objects within a closed-loop control system. Our framework leverages a vision-language model for natural language processing and real-time object segmentation, allowing for the seamless integration of 6D object pose localization and receding-horizon trajectory optimization to achieve precise and smooth robotic movements. The intensive experimental results highlight the robustness and efficiency of our framework, achieving a 92% success rate and superior inference speeds compared to state-of-the-art methods. Furthermore, the object localization module demonstrated high reliability with minimal positional and orientation errors, particularly in dynamic environments requiring real-time replanning. Future work will focus on further improving 6D pose estimation by continuously localizing object segmentation to prevent tracking loss and detecting affordance-based grasp points for objects through a language-driven approach.